OpenStudy (anonymous):

Does the sequence converge, or diverge? (-1)^(n-1)/sqrt(n)

OpenStudy (anonymous):

From 1 to infinity.

OpenStudy (anonymous):

oop it is a SEQUENCE right, not a SERIES

OpenStudy (anonymous):

\[a_n=\frac{(-1)^{n-1}}{\sqrt{n}}\]

OpenStudy (anonymous):

as \(n\to \infty\) it should be pretty clear that \(a_n\to 0\)

OpenStudy (anonymous):

How though? That's what I don't understand..The steps to get that answer. @satellite73

OpenStudy (anonymous):

I would suggest showing that the sequence is convergent by showing it is both monotonic and bounded. Since you have an alternating series, the monotonicity part might be tricky. As a hint, try working with the positive and negative terms separately.

OpenStudy (anonymous):

@SithsAndGiggles any idea to start it?

OpenStudy (anonymous):

\[\{a_n\}=\left\{\frac{(-1)^{n-1}}{\sqrt{n}}\right\}_{n=1}^\infty=\left\{1,-\frac{1}{\sqrt2},\frac{1}{\sqrt3},-\frac{1}{2},\cdots\right\}\] Consider \(\{b_n\}=\left\{\dfrac{(-1)^{2k-1}}{\sqrt{2k}}\right\}_{k=1}^\infty\) and \(\{c_n\}=\left\{\dfrac{(-1)^{2k}}{\sqrt{2k+1}}\right\}_{k=0}^\infty\). Notice that \(b_n\) is the subsequence of \(a_n\) consisting of the even terms (\(a_2,a_4,\cdots\)) and \(c_n\) is the subsequence of the odd terms (\(a_1,a_3,\cdots\)). Showing that \(b_n\) and \(c_n\) converge should be enough to show that \(a_n\) converges. Alternatively, you can show the following conditions hold: (1) the magnitude of each term is decreasing, meaning \(|a_n|\ge|a_{n+1}|\), and (2) the terms converge to 0. I think the second method would be a more valid approach. The first is a bit more hand-wavy.

OpenStudy (anonymous):

So I should just prove that An converges?

OpenStudy (anonymous):

\(a_n\) is what you're given, so yes, that's what you want to show convergence for. Show that the conditions are met and you're done.